[figure: 9.7 cm]

NONCOLLINEAR SPIN CONFIGURATIONS IN Fe/Si MULTILAYERS

A.B. Chizhik^{1}, K. Fronc^{2}, S.L. Gnatchenko^{1}, D.N. Merenkov^{1}, R. Żuberek^{2} 1 Institute for Low Temperature Physics & Engineering, National Academy of Sciences

of Ukraine, 310164 Kharkov, Ukraine 2 Institute of Physics, Polish Academy of Sciences, 02-668 Warsaw, Poland

Formation of noncollinear spin configurations has been investigated in (Fe30Å/Si15Å)·11 multilayered films, which were prepared by sequential dc sputtering. A single-crystal GaAs plate of (001) orientation was used as substrate. The experiments have been performed using magneto-optical longitudinal Kerr effect at room temperature. The magnetic field was applied in the film plane. The field dependencies of the Kerr rotation had parts with slightly varying rotation. For different directions of magnetic field in the film plane the Kerr curves had different value of F_{0}/F_{s} ratio (F_{s} and F_{0} are the Kerr rotation in saturation and in zero magnetic field, respectively). The presence of cubic anisotropy and weak uniaxial anisotropy in the film plane has been demonstrated by the ferromagnetic resonance measurements. The experimental results were analyzed using the theory, which takes into account the competition between the bilinear exchange term, biquadratic exchange term and cubic magnetic anisotropy term. As the analysis shows, the observed features are related to the field-induced phase transition to stable noncollinear structure, which exists as metastable state in the absence of magnetic field.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : A.B. Chizhik

Address for correspondence : Institute for Low Temperature Physics & Engineering, 47 Lenin Ave., 310164 Kharkov, UKRAINE

Email address : chizhik@ilt.kharkov.ua